OpenStudy (anonymous):

Find the value of each expression. 9P3

OpenStudy (anonymous):

9P3?

OpenStudy (shubhamsrg):

nPr = n! /(n-r)! whats the problem ?

OpenStudy (anonymous):

\[_{9}P _{3}\]

OpenStudy (anonymous):

ahh .. nothing . .hehe .. :D don't mind it .. :D

OpenStudy (anonymous):

Can you show me how to find the answer.

OpenStudy (hesan):

by calculator u can find answer... :)

OpenStudy (anonymous):

\[\left\{ 27P \right\}\]

OpenStudy (anonymous):

9P3 = 9! / (9-3)! = 9·8·7·6! / 6! = 9·8·7 =

OpenStudy (anonymous):

\[\frac{ 362880 }{ 720 }=504\]

OpenStudy (anonymous):

Thank you for helping me I couldn't figure out what I was looking for in my book.

OpenStudy (anonymous):

I message you if got other question just post. Welcome to OS and give medal to the one you like.